# Dimitrios M. Thilikos ${}^{,}$

### v3

@@ 5.1 A local structure theorem for excluding Dyck-grids
In case $c=2n\neq 0,$ let $h^{\prime}\coloneqq h+n-1$ and $c^{\prime}\coloneqq 2$, similarly, if $c=0$ we set $h^{\prime}\coloneqq h$ and $c^{\prime}\coloneqq c=0$. Moreover, if $c=2n+1,$ let $h^{\prime}\coloneqq h+n$ and $c^{\prime}\coloneqq 1.$ By Subsection 5.1 we find $D\coloneqq\mathscr{D}^{(h^{\prime},c^{\prime})}_{12k}$ as a minor of $G.$
What follows is an analysis on the numbers $h^{\prime},c^{\prime},c,$ and $h$ in relation to $g_{1}$ and $g_{2}.$
@@
Suppose $h^{\prime}\geq h_{1},$ then $D$ contains $\mathscr{D}_{k}^{(h_{1},0)}$ as a minor, and we are done. So we may assume $h^{\prime}<h_{1}$ in the following. This settles the cases where $c=0$ or $c_{2}=0$ since both situations imply that $2h^{\prime}+c^{\prime}\leq 2h<g_{1}.$
From now on we may assume $c,c_{2}\neq 0$ which implies that $c^{\prime},c_{2}\in[2].$
@@
Let us first assume that $c^{\prime}=1$ and $2h^{\prime}+c^{\prime}\geq 2h_{2}+c_{2}.$ If $c_{2}=1$ we immediately obtain $h^{\prime}\geq h_{2}.$ Hence, we may find $\mathscr{D}^{(h_{2},c_{2})}_{k}$ as a minor in $D$ as desired. So we may assume $c_{2}=2.$ In this case, however, since the two sums have different parity we obtain $h^{\prime}>h_{2}.$ We now use Subsection 2.3 to obtain the surface grid $D^{\prime}$ of order $k$ with $h^{\prime}-1$ handles and $c^{\prime}+2$ crosscaps as a minor within $D.$ As $h^{\prime}-1\geq h_{2}$ and certainly $c^{\prime}+2>c_{2}$ we find $\mathscr{D}^{(h_{2},c_{2})}_{k}$ as a minor in $D^{\prime}.$
@@
For the final case we assume $c^{\prime}=2$ and again $2h^{\prime}+c^{\prime}\geq 2h_{2}+c_{2}.$ Since this means $c^{\prime}\geq 2\geq c_{2}\geq 1$ we also obtain $h^{\prime}\geq h_{2}$ which immediately implies the presence of a $\mathscr{D}^{(h_{2},c_{2})}_{k}$-minor in $D$ as desired and our proof is complete. ∎
5.2 Proof of Theorem 5.1
@@
With Subsection 5.1 at hand our proof of Theorem 5.1 becomes an analogue of the proof for Theorem 2.11 from [32]. However, we split the proof in half to make an intermediate structure accessible for future applications. The reason for this is the requirement of a local structure theorem to produce a $\Sigma$-decomposition which is $W$-central with respect to a large wall. The way $W$-centrality is achieved in [32] is by removing additional vertex sets from the vortices which potentially damages the surrounding nests. Some applications might need access to the nests around vortices after the fact, and thus we include an intermediate step here.
The main reason most of the structural theory talks about walls rather than grids is that walls are subcubic graphs and as such, if they are contained as minors in a graph, one can find a subdivision as a subgraph. For a similar reason we define Dyck-walls. Let $c\in[0,2]$, $h$, and $t$ be non-negative integers. The elementary $(h,c;t)$-Dyck-wall is obtained from $\mathcal{D}^{h,c}_{2t}$ by deleting the cycles $C_{t+1},\dots,C_{2t}$ and by deleting the edge $v^{i}_{j}v^{i+1}_{j}$ for every odd $i\in[t-1]$ and every odd $j\in[8t]$ and for every even $i\in[t-1]$ and even $j\in[8t]$. Moreover, for each handle and crosscap that was added to create $\mathcal{D}^{h,c}_{2t}$ delete every edge starting on $v^{1}_{\ell}$ where $\ell$ is even. That is, we delete every second edge. An $(h,c;t)$-Dyck-wall is a subdivision of the elementary $(h,c;t)$-Dyck-wall.
Consider an embedding of a $(h,c;t)$-Dyck-wall $D$ in a surface $\Sigma$ of Euler-genus $2h+c$. Then $C_{t}$ bounds a disk in which no vertex of $D$ is drawn. We call this disk the simple face of $D$. Similarly, there exists a unique face which is incident to $C_{1}$ and every handle and every crosscap of $D$. We call this face the exceptional face of $D$.
@@
We call a graph $H$ that appears as the torso of a node in a tree-decomposition of $G,$ as in Subsection 5.2, a quasi-$4$-connencted component of $G$.
@@
Let $h_{1},h_{2},c_{2}$ be non-negative integers where $c_{2}\in[0,2]$ and $0\leq h_{2}<h_{1}.$ We set
$g_{1}$$\coloneqq$ $2h_{1},$
@@
Let also $k$, $r$, q be non-negative integers, where, $r\geq 16\cdot 12(162^{2g}+4)k$ and $q\geq 6k$, and set
$t$$\coloneqq$ $12(162^{2g}+4)k,$
@@
    $G$ contains $\mathscr{D}^{(h_{1},0)}_{k}$ or, in case $c_{2}\neq 0,$ $\mathscr{D}^{(h_{2},c_{2})}_{k}$ as a minor, or
@@
    • a surface $\Sigma\coloneqq\Sigma^{(h,c)}$ where
      • $2h<g_{1}$,
      • if $c\neq 0$ and $c_{2}\neq 0$ then $2h+c<g_{2}$, and
      • if $c\neq 0$ and $c_{2}=0$ then $2h+c\leq g_{1}$,
@@
    • a $\Sigma$-decomposition $\delta=(\Gamma,\mathcal{D})$ of $G-A$ with the following properties:
@@
        if we define $c_{0}$ and $h_{0}$ as follows:
        • if $c=0$, let $c_{0}\coloneqq 0$ and $h_{0}\coloneqq h$,
        • if $c>0$ and $c+2h$ is even, let $c_{0}\coloneqq 2$ and $h_{0}\coloneqq\frac{c+2h}{2}-1$, and
        • otherwise, let $c_{0}\coloneqq 1$ and $h_{0}\coloneqq\frac{c+2h-1}{2}$,
        then there exists an $(h_{0},c_{0};q)$-Dyck-wall $D$ in $G-A$ such that $D$ is grounded in $\delta$
@@
        there exists an $r$-subwall $W^{\prime}$ of $W$ such that $W^{\prime}$ is grounded in $\delta,$ and $W^{\prime}$ is drawn in the interior of the disk defined by the trace of the boundary of the simple face of $D$,
@@
        for each vortex cell $v\in C(\delta)$ there exists a nest $\mathcal{C}_{v}$ of order $M$ in $\delta$ around the unique disk $\Delta\in\mathcal{D}$ corresponding to $v,$
@@
        let $X$ be the set of all vertices of $G$ drawn in the interior of cells of $\delta$ and let $G_{\delta}$ be the torso of the set $V(G)\setminus(X\cup A)$ in $G-A$, then $G_{\delta}$ is a quasi-$4$-connected component of $G-A$, and
@@
By our choice of parameters we may immediatelly call upon Subsection 5.1. The first possible outcome of Subsection 5.1 yields a minor model of $\mathscr{D}^{(h_{1},0)}_{k}$ or, in case $c_{2}\neq 0,$ of $\mathscr{D}^{(h_{2},c_{2})}_{k}.$
The second outcome provides us with integers $c,h\geq 0$ with $2h<g_{1},$ $2h+c<g_{2}$ if $c\neq 0$ and $c_{2}\neq 0,$ and $2h+c\leq g_{1}$ if $c\neq 0$ and $c_{2}=0,$ a set $A\subseteq V(G)$ of size at most $\mu,$ a surface $\Sigma$ of Euler-genus $c+2h,$ which can be obtained from the sphere by adding $h$ handles and $c$ crosscaps, and a $\Sigma$-configuration $(W_{1},W_{0},\delta,o_{0},o_{1},\Delta_{0},\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{P}_{1},\mathcal{P}_{2},\gamma)$ of $G-A$ with parameters $(r,M,2M,M,M,c,h)$.
Additionally, we have that the vortex society of $o_{1}$ has a rendition $\rho_{1}=(\Gamma_{1},\mathcal{D}_{1})$ of breadth at most $2t^{2}$ and depth at most $\mu$ such that
  • for each vortex cell $c\in C(\rho_{1})$ there exists a nest $\mathcal{C}_{c}$ of order $M$ in $\rho_{1}$ around the unique disk $\Delta\in\mathcal{D}$ corresponding to $c,$ and
@@
We extend $\delta$ to a $\Sigma$-decomposition $\delta^{\prime}=(\Gamma^{\prime},\mathcal{D})$ of $G-A$ by fixing a vortex free rendition of the vortex society of $o_{0}$ such that the wall $W_{0}$ is flat in the rendition and fixing a rendition of breath at most $2t^{2}$ and depth at most $\mu$ of the vortex society of $o_{1}.$ Thus, $\delta^{\prime}$ has breath at most $2t^{2}$ and depth at most $\mu.$
@@
Next, Subsection 5.1 provides us with a minor-model $D^{\prime}$ of $\mathscr{D}^{h_{0},c_{0}}_{2q}$ which is grounded in $\delta^{\prime}$. Notice that $D^{\prime}$ contains an $(h_{0},c_{0};q)$-Dyck-wall $D$ as a subgraph which still is grounded in $\delta^{\prime}$. In particular, $D^{\prime}$ is also grounded in $\delta$, implying that $D$ is as well. From the construction of $D^{\prime}$ from the mixed surface grid in Subsection 5.1 it follows that $o_{0}$ sits in the disk bounded by the trace of the simple face of $D$ while $o_{1}$ belongs to the disk bounded by the exceptional face of $D$.
@@
We now start to further modify $\delta^{\prime}$ by iterating the following operation exhaustively. Let $(Y_{1},Y_{2})$ be a separation of order at most three in $G-A$ such that $Y_{1}\cap Y_{2}$ consists of nodes of $\delta^{\prime}$ and assume that $(Y_{1},Y_{2})$ does not cross any separation induced by a non-vortex cell of $\delta^{\prime}$. That is, we assume that there does not exist a non-vertex cell $x\in C(\delta^{\prime})$ such that $V(\sigma^{\prime}(x))\cap(Y_{1}\setminus Y_{2})\neq\emptyset$ and $V(\sigma^{\prime}(x))\cap(Y_{2}\setminus Y_{1})\neq\emptyset$. By the submodularity of separations we may always find such a separation $(Y_{1},Y_{2})$ if some separation of order at most three with only nodes in its separator exists. We start with separations of order one, then consider separations of order two, and finally separations of order three. Without loss of generality we may assume that $Y_{1}$ contains a row and a column of $W_{0}$. Moreover, we assume that $Y_{2}$ is maximal with respect to set containment. This implies that $Y_{1}\setminus Y_{2}$ is connected. Indeed, it follows from the definition of $\Sigma$-configurations and the $3$-connectivity of Dyck-grids that also at most one branch vertex of $D$ may be contained in $Y_{2}\setminus Y_{1}$. Now suppose $Y_{2}\setminus Y_{1}$ contains a node of $\delta^{\prime}$. Let $C_{(Y_{1},Y_{2})}$ be the collection of all cells $x$ of $\delta^{\prime}$ such that $V(\sigma^{\prime}(x))\cap(Y_{2}\setminus Y_{1})\neq\emptyset$. It follows that $V(\sigma^{\prime}(x))\subseteq Y_{2}$ for all $x\in C_{(Y_{1},Y_{2})}$ by our choice of $(Y_{1},Y_{2})$. We may now define a new cell $c_{(Y_{1},Y_{2})}$ replacing all cells in $C_{(Y_{1},Y_{2})}$ as follows. Let $\Delta_{(Y_{1},Y_{2})}$ be the disk which is the closure of $\bigcup_{c\in C_{(Y_{1},Y_{2})}}c$. Now let $c_{(Y_{1},Y_{2})}$ be the cell obtained by removing the points $\pi^{\prime-1}(y)$, $y\in Y_{1}\cap Y_{2}$, from $\Delta_{(Y_{1},Y_{2})}$, and notice that all vertices of $Y_{2}\setminus Y_{1}$ are drawn into the interior of $c_{(Y_{1},Y_{2})}$. Then forget all cells in $C_{(Y_{1},Y_{2})}$. Let $\delta^{\prime\prime}=(\Gamma^{\prime\prime},\mathcal{D}^{\prime\prime})$ be the $\Sigma$-decomposition obtained once the above operation cannot be applied any more. Notice that $D$ is still grounded in $\delta^{\prime\prime}$ and so is $W_{0}$. The same is true for all cycles from the nests of the vortices of $\delta^{\prime\prime}$. This implies that the torso $G_{\delta^{\prime\prime}}$ contains a Dyck-wall $D^{\prime\prime}$ of the same order as $D$ which can be obtained by contracting some edges of $D$ and is quasi-$4$-connected.
@@
The proof of Subsection 5.2 can be turned into an algorithm that given a graph $G$, $k,h_{1},h_{2},c_{2},r,M$ and an $R$-wall as above, computes the structures implied by Subsection 5.2 in time
$\mathcal{O}(M^{2}\cdot 2^{\mathsf{poly}(k\cdot 2^{\max\{2h_{1},2h_{2}+c_{2}\}})}|V(G)|^{2}+|V(G)|^{3}).$
@@
We may call upon Subsection 5.2 with the numbers $k,h_{1},h_{2},c_{2},$ $r^{\prime},s_{2},M$ and $\mu$ and the wall $W.$ In case we find $\mathscr{D}^{(h_{1},0)}_{k}$ or $\mathscr{D}^{(h_{2},c_{2})}_{k}$ as a minor we are done, so we may assume the final outcome of Subsection 5.2 to hold.
This means we find a set $A\subseteq V(G)$ of size at most $\mu$ and $\Sigma$-decomposition $\delta=(\Gamma,\mathcal{D})$ of $G-A$ satisfying properties a) to h) from Subsection 5.2. What remains is to ensure that $\delta$ is $W$-central.

### v4

@@ 6.1 Extracting a Dyck-wall
In case $\mathsf{c}=2n\neq 0,$ let $\mathsf{h}^{\prime}\coloneqq\mathsf{h}+n-1$ and $\mathsf{c}^{\prime}\coloneqq 2,$ similarly, if $\mathsf{c}=0$ we set $\mathsf{h}^{\prime}\coloneqq\mathsf{h}$ and $\mathsf{c}^{\prime}\coloneqq\mathsf{c}=0.$ Moreover, if $\mathsf{c}=2n+1,$ let $\mathsf{h}^{\prime}\coloneqq\mathsf{h}+n$ and $\mathsf{c}^{\prime}\coloneqq 1.$ By Subsection 6.1 we find $D\coloneqq\mathscr{D}^{(\mathsf{h}^{\prime},\mathsf{c}^{\prime})}_{k}$ as a minor of $G.$
Observe that, by Subsection 3.3, if $K_{t}$ embeds in $\Sigma^{(\mathsf{h},\mathsf{c})},$ then we find our $K_{t}$-minor grasped by $W$ within $D.$ Moreover, if $\mathscr{D}^{(\mathsf{h}_{1},0)}_{k}$ embeds in $\Sigma^{(\mathsf{h},\mathsf{c})},$ then $D$ contains $\mathscr{D}^{(\mathsf{h}_{1},0)}_{k}$ as a minor in a way that lets their tangles agree. So the case where $\mathsf{c}^{\prime}\neq 0$ but $\mathsf{c}_{2}=0$ is closed.
What follows is an analysis on the numbers $\mathsf{h}^{\prime},\mathsf{c}^{\prime},\mathsf{c},$ and $\mathsf{h}$ in relation to $g_{1}=2\mathsf{h}_{1}$ and $g_{2}=2\mathsf{h}_{2}+\mathsf{c}_{2}.$
@@
Suppose $\mathsf{h}^{\prime}\geq\mathsf{h}_{1},$ then $D$ contains $\mathscr{D}_{k}^{(\mathsf{h}_{1},0)}$ as a minor, and we are done. So we may assume $\mathsf{h}^{\prime}<\mathsf{h}_{1}$ in the following. This settles the cases where $\mathsf{c}=0$ or $\mathsf{c}_{2}=0$ since both situations imply that $2\mathsf{h}^{\prime}+\mathsf{c}^{\prime}\leq 2\mathsf{h}<g_{1}.$
From now on we may assume $\mathsf{c},\mathsf{c}_{2}\neq 0$ which implies that $\mathsf{c}^{\prime},\mathsf{c}_{2}\in[2].$
@@
Let us first assume that $\mathsf{c}^{\prime}=1$ and $2\mathsf{h}^{\prime}+\mathsf{c}^{\prime}\geq 2\mathsf{h}_{2}+\mathsf{c}_{2}.$ If $\mathsf{c}_{2}=1$ we immediately obtain $\mathsf{h}^{\prime}\geq\mathsf{h}_{2}.$ Hence, we may find $\mathscr{D}^{(\mathsf{h}_{2},\mathsf{c}_{2})}_{k}$ as a minor in $D$ as desired. So we may further assume that $\mathsf{c}_{2}=2.$ Since the two sums have different parity, we obtain $\mathsf{h}^{\prime}>\mathsf{h}_{2}.$ We now use Subsection 3.1 to obtain the surface grid $D^{\prime}$ of order $k$ with $\mathsf{h}^{\prime}-1$ handles and $\mathsf{c}^{\prime}+2$ crosscaps as a minor within $D.$ As $\mathsf{h}^{\prime}-1\geq\mathsf{h}_{2}$ and certainly $\mathsf{c}^{\prime}+2>\mathsf{c}_{2}$ we find $\mathscr{D}^{(\mathsf{h}_{2},\mathsf{c}_{2})}_{k}$ as a minor in $D^{\prime}$ as desired.
@@
For the final case we assume $\mathsf{c}^{\prime}=2$ and again $2\mathsf{h}^{\prime}+\mathsf{c}^{\prime}\geq 2\mathsf{h}_{2}+\mathsf{c}_{2}.$ Since this means $\mathsf{c}^{\prime}\geq 2\geq\mathsf{c}_{2}\geq 1$ we also obtain $\mathsf{h}^{\prime}\geq\mathsf{h}_{2}$ which immediately implies the presence of a $\mathscr{D}^{(\mathsf{h}_{2},\mathsf{c}_{2})}_{k}$-minor in $D$ as desired.
What is left is to ensure that $\delta$ is indeed $W$-central. ∎
6.2 The proof of Theorem 6.1
@@
With Subsection 6.1 at hand our proof of Theorem 6.1 becomes an analogue of the proof for Theorem 2.11 from [17]. However, we split the proof in half to make an intermediate structure accessible for future applications. The reason for this is that we require our final $\Sigma$-decomposition to be $W$-central with respect to a large wall $W.$ The way $W$-centrality is achieved in [17] is by removing additional vertex sets from the vortices which potentially damages the surrounding nests. Some applications might need access to the nests around vortices after the fact, and thus we include an intermediate step here.
@@
We call a graph $H$ that appears as the torso of a node in a tree-decomposition of $G,$ as in Subsection 6.2, a quasi-$4$-connected component of $G.$
@@
Let $\mathsf{h}_{1},\mathsf{h}_{2},\mathsf{c}_{2}$ be non-negative integers where $\mathsf{c}_{2}\in[0,2]$ and $0\leq\mathsf{h}_{2}<\mathsf{h}_{1}.$ We set
$g_{1}$$\coloneqq$ $2\mathsf{h}_{1},$
@@
Let also $t\geq 1,$ $k,$ $r,$ and $q$ be non-negative integers, where, $r\geq 16\cdot(162^{2g}+4)k$ and $q\geq\max\{k,t\},$ and set
@@
    $H$ contains a $K_{t}$-minor grasped by $W,$
@@
    $G$ contains an expansion $D$ of (a) $\mathscr{D}^{(\mathsf{h}_{1},0)}_{k}$ or (b) $\mathscr{D}^{(\mathsf{h}_{2},\mathsf{c}_{2})}_{k},$ where outcome (b) is allowed if and only if $\mathsf{c}_{2}\neq 0,$ such that $\mathcal{T}_{D}\subseteq\mathcal{T}_{W},$ or
  3. 3.
@@
    • a surface $\Sigma\coloneqq\Sigma^{(\mathsf{h},\mathsf{c})}$ where $2\mathsf{h}<2\mathsf{h}_{1}$ in case $c=0,$ and $2\mathsf{h}+\mathsf{c}<2\mathsf{h}_{2}+\mathsf{c}_{2}$ if $\mathsf{c}_{2}\neq 0.$
@@
    • a $\Sigma$-decomposition $\delta=(\Gamma,\mathscr{D})$ of $G-A$ with the following properties:
@@
        if we define $\mathsf{c}_{0}$ and $\mathsf{h}_{0}$ as follows:
        • if $\mathsf{c}=0,$ let $\mathsf{c}_{0}\coloneqq 0$ and $\mathsf{h}_{0}\coloneqq\mathsf{h},$
        • if $\mathsf{c}>0$ and $\mathsf{c}+2\mathsf{h}$ is even, let $\mathsf{c}_{0}\coloneqq 2$ and $\mathsf{h}_{0}\coloneqq\frac{\mathsf{c}+2\mathsf{h}}{2}-1,$ and
        • otherwise, let $\mathsf{c}_{0}\coloneqq 1$ and $\mathsf{h}_{0}\coloneqq\frac{\mathsf{c}+2\mathsf{h}-1}{2},$
        then there exists an $(\mathsf{h}_{0},\mathsf{c}_{0};q)$-Dyck-wall $D$ in $G-A$ such that $D$ is grounded in $\delta$ and $\mathcal{T}_{D}\subseteq\mathcal{T}_{W}$
@@
        there exists an $r$-subwall $W^{\prime}$ of $W$ such that $W^{\prime}$ is flat in $\delta,$ and $W^{\prime}$ is drawn in the interior of the disk defined by the trace of the boundary of the simple face of $D,$
@@
        for each vortex cell $v\in C(\delta)$ there exists a nest $\mathcal{C}_{v}$ of order $M$ in $\delta$ around the unique disk $\Delta\in\mathscr{D}$ corresponding to $v,$
@@
        let $X$ be the set of all vertices of $G$ drawn in the interior of cells of $\delta$ and let $G_{\delta}$ be the torso of the set $V(G)\setminus(X\cup A)$ in $G-A,$ then $G_{\delta}$ is a quasi-$4$-connected component of $G-A,$ and margin: SED> $G_{\delta}$
@@
By our choice of parameters we may immediately call upon Subsection 6.1. The first possible outcome of Subsection 6.1 yields a $K_{t}$-minor grasped by $W$ as desired. The second outcome of Subsection 6.1 provides us with a minor model of $\mathscr{D}^{(\mathsf{h}_{1},0)}_{k}$ or, in case ${\sf c}_{2}\neq 0,$ of $\mathscr{D}^{(\mathsf{h}_{2},\mathsf{c}_{2})}_{k}.$
The third outcome provides us with integers $\mathsf{c},\mathsf{h}\geq 0$ and s surface $\Sigma=\Sigma^{(\mathsf{h},\mathsf{c})}$ where none of the graphs $K_{t},$ $\mathscr{D}^{(\mathsf{h}_{1},0)}_{k},$ and, if ${\sf c}_{2}\neq 0,$ $\mathscr{D}^{(\mathsf{h}_{2},\mathsf{c}_{2})}_{k}$ embeds. Moreover, we obtain a set $A\subseteq V(G)$ of size at most $\mu$ and a $\Sigma$-configuration $(W_{1},W_{0},\delta,o_{0},o_{1},\Delta_{0},\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{P}_{1},\mathcal{P}_{2},\gamma)$ of $G-A$ with parameters $(r,M,2M,M,M,\mathsf{c},\mathsf{h}).$
Additionally, we have that the vortex society of $o_{1}$ has a rendition $\rho_{1}=(\Gamma_{1},\mathscr{D}_{1})$ of breadth at most $2t^{2}$ and depth at most $\mu$ such that
  • for each vortex cell $c\in C(\rho_{1})$ there exists a nest $\mathcal{C}_{c}$ of order $M$ in $\rho_{1}$ around the unique disk $\Delta\in\mathscr{D}$ corresponding to $c,$ and
@@
We extend $\delta$ to a $\Sigma$-decomposition $\delta^{\prime}=(\Gamma^{\prime},\mathscr{D})$ of $G-A$ by fixing a vortex free rendition of the vortex society of $o_{0}$ such that the wall $W_{0}$ is flat in the rendition and fixing a rendition of breath at most $2t^{2}$ and depth at most $\mu$ of the vortex society of $o_{1}.$ Thus, $\delta^{\prime}$ has breath at most $2t^{2}$ and depth at most $\mu.$
@@
Next, Subsection 6.1 provides us with a minor-model $D^{\prime}$ of $\mathscr{D}^{(\mathsf{h}_{0},\mathsf{c}_{0})}_{2q}$ which is grounded in $\delta^{\prime}.$ Notice that $D^{\prime}$ contains an $(\mathsf{h}_{0},\mathsf{c}_{0};q)$-Dyck-wall $D$ as a subgraph which still is grounded in $\delta^{\prime}.$ In particular, $D^{\prime}$ is also grounded in $\delta,$ implying that $D$ is as well. From the construction of $D^{\prime}$ from the mixed surface grid in Subsection 6.1 it follows that $o_{0}$ sits in the disk bounded by the trace of the simple face of $D$ while $o_{1}$ belongs to the disk bounded by the exceptional face of $D.$
We now start to further modify $\delta^{\prime}$ by iterating the following operation exhaustively. Let $(Y_{1},Y_{2})$ be a separation of order at most three in $G-A$ such that $Y_{1}\cap Y_{2}$ consists of nodes of $\delta^{\prime}$ and assume that $(Y_{1},Y_{2})$ does not cross any separation induced by a non-vortex cell of $\delta^{\prime}.$ That is, we assume that there does not exist a non-vertex cell $x\in C(\delta^{\prime})$ such that
$V(\sigma^{\prime}(x))\cap(Y_{1}\setminus Y_{2})\neq\emptyset\text{\leavevmode\nobreak\ and\leavevmode\nobreak\ }V(\sigma^{\prime}(x))\cap(Y_{2}\setminus Y_{1})\neq\emptyset.$
@@
By the submodularity of separations we may always find such a separation $(Y_{1},Y_{2})$ if some separation of order at most three with only nodes in its separator exists. We start with separations of order one, then consider separations of order two, and finally separations of order three. Without loss of generality we may assume that $Y_{1}$ contains a row and a column of $W_{0}.$ Moreover, we assume that $Y_{2}$ is maximal with respect to set containment. This implies that $Y_{1}\setminus Y_{2}$ is connected. Indeed, it follows from the definition of $\Sigma$-configurations and the $3$-connectivity of Dyck-grids that also at most one branch vertex of $D$ may be contained in $Y_{2}\setminus Y_{1}.$ Now suppose $Y_{2}\setminus Y_{1}$ contains a node of $\delta^{\prime}.$ Let $C_{(Y_{1},Y_{2})}$ be the collection of all cells $x$ of $\delta^{\prime}$ such that $V(\Sigma^{\prime}(x))\cap(Y_{2}\setminus Y_{1})\neq\emptyset.$ It follows that $V(\sigma^{\prime}(x))\subseteq Y_{2}$ for all $x\in C_{(Y_{1},Y_{2})}$ by our choice of $(Y_{1},Y_{2}).$ We may now define a new cell $c_{(Y_{1},Y_{2})}$ replacing all cells in $C_{(Y_{1},Y_{2})}$ as follows. Let $\Delta_{(Y_{1},Y_{2})}$ be the disk which is the closure of $\bigcup_{c\in C_{(Y_{1},Y_{2})}}c.$ Now let $c_{(Y_{1},Y_{2})}$ be the cell obtained by removing the points $\pi^{\prime-1}(y),$ $y\in Y_{1}\cap Y_{2},$ from $\Delta_{(Y_{1},Y_{2})},$ and notice that all vertices of $Y_{2}\setminus Y_{1}$ are drawn into the interior of $c_{(Y_{1},Y_{2})}.$ Then forget all cells in $C_{(Y_{1},Y_{2})}.$ Let $\delta^{\prime\prime}=(\Gamma^{\prime\prime},\mathscr{D}^{\prime\prime})$ be the $\Sigma$-decomposition obtained once the above operation cannot be applied any more. Notice that $D$ is still grounded in $\delta^{\prime\prime}$ and so is $W_{0}.$ The same is true for all cycles from the nests of the vortices of $\delta^{\prime\prime}.$ This implies that the torso $G_{\delta^{\prime\prime}}$ contains a Dyck-wall $D^{\prime\prime}$ of the same order as $D$ which can be obtained by contracting some edges of $D$ and is quasi-$4$-connected.
@@
Deriving Theorem 1.1.
With regards to improving the bound on the Euler-genus from [17] let us fix some graph $H.$ We set $g_{1}$ to be the smallest Euler genus of an orientable surface where $H$ embeds and $g_{2}$ to be the smallest Euler genus of a non-orientable surface where $H$ embeds. We then set $t\coloneqq|V(H)|.$ Then, one needs to set $k$ to be $\Theta\big{(}162^{2\max\{g_{1},g_{2}\}}(\max\{g_{1},g_{2}\})+t^{2}\big{)}.$ This ensures that Subsection 3.3 witnesses the presence of an $H$-minor, even if the surface resulting from Section 5 has an Euler genus below the bound guaranteed there. This concludes the proof of Theorem 1.1.
The proof of Subsection 6.2 can be turned into an algorithm that given a graph $G,$ $k,\mathsf{h}_{1},\mathsf{h}_{2},\mathsf{c}_{2},r,M$ and an $R$-wall as above, computes the structures implied by Subsection 6.2 in time
[…]
@@
We may call upon Subsection 6.2 with the numbers $k,\mathsf{h}_{1},\mathsf{h}_{2},\mathsf{c}_{2},$ $r^{\prime},s_{2},M$ and $\mu$ and the wall $W.$ In case we find $K_{t},$ $\mathscr{D}^{(\mathsf{h}_{1},0)}_{k},$ or $\mathscr{D}^{(\mathsf{h}_{2},\mathsf{c}_{2})}_{k}$ as a minor we are done. This is particularly true by our choice of $t$ as in case we find the $K_{t}$-minor, it will contain both $\mathscr{D}^{(\mathsf{h}_{1},0)}_{k}$ and $\mathscr{D}^{(\mathsf{h}_{2},\mathsf{c}_{2})}_{k}$ as minors. So we may assume the final outcome of Subsection 6.2 to hold.
This means we find a set $A\subseteq V(G)$ of size at most $\mu$ and $\Sigma$-decomposition $\delta=(\Gamma,\mathscr{D})$ of $G-A$ satisfying properties a) to h) from Subsection 6.2. What remains is to ensure that $\delta$ is $W$-central.
@@
7 From local to global structure theorems
Finally, we have everything in place to prove a local version of our, so far local, structural theorems. To close the gap in the Euler-genus left by Kawarabayashi, Thomas, and Wollan [17] we would need to provide a global version of Subsection 6.2. However, the main goal of this paper is to understand the resulting structure when the universal graphs of some orientable and some non-orientable surface are excluded. Since the general strategy of the proof, we follow the proof of [17], an earlier version of this had appeared in [22], stays the same for all such “globalisation”-type arguments, we only present here the global version of Theorem 6.1. Proving a global version of Subsection 6.2 follows along the same line of arguments.
Let $k,$ $\mathsf{h}_{1},$ $\mathsf{h}_{2},$ $\mathsf{c}_{2},$ and $r$ be non-negative integers with $\mathsf{c}_{2}\in[0,2]$ and $0\leq\mathsf{h}_{2}<\mathsf{h}_{1}.$ Let $g_{1}\coloneqq 2\mathsf{h}_{1},$ $g_{2}\coloneqq 2\mathsf{h}_{2}+\mathsf{c}_{2}$ if $\mathsf{c}_{2}\neq 0$ and $g_{2}\coloneqq 0$ otherwise. Moreover, let $g\coloneqq\max\{g_{1},g_{2}\}$ and fix $t\coloneqq(162^{2g}+4)k.$ Recall the definitions of the functions $\mathbf{a}_{g_{1},g_{2}}=t^{10^{7}t^{26}},$ $\mathbf{v}_{\mathsf{b},g_{1},g_{2}}=2t^{2},$ and $\mathbf{v}_{\mathsf{d},g_{1},g_{2}}=t^{10^{7}t^{26}}$ from Theorem 6.1. We define the following function. Let $c_{1}$ be the constant from Theorem 4.1. We set
$\mathbf{a}(g_{1},g_{2},k,r)$ $\coloneqq 36c_{1}\Big{(}49152t^{24}(r+16t)+\left\lceil\frac{5}{2}t^{10^{7}t^{26}}\right\rceil\Big{)}^{20}+4\text{, with}.$